OpenStudy (anonymous):

If f(x)=x^2 +x , find f(x-h)

OpenStudy (anonymous):

f(x-h) is the value of function f when x = x-h So f(x-h) = (x-h)^2 + (x-h) You may want to simplify the expression.

OpenStudy (anonymous):

Ok, thanks!

OpenStudy (anonymous):

welcome.